# Triangular solutions to a Riemann-Hilbert problem from superelliptic curves

### v1

@@ 1.  Introduction
Once a fundamental solution to (1) is found, one may obtain its monodromy matrices. The inverse problem, that of reconstructing a function $\Phi(z)$ from its monodromies, is referred to as a Riemann-Hilbert or inverse monodromy problem. Alternatively, it may be reformulated as a problem of finding two matrix functions, each of which is holomorphic either inside or outside of a closed contour passing trough the points $a_{1},\dots,a_{N}$ in the complex plane, such that the two functions are related by the multiplication with a “jump” matrix on the contour. The monodromy matrices associated with the loops around $a_{i}$ may be combined to obtain this jump matrix [4, 17]. The term Riemann-Hilbert problem refers also to the 21st problem from Hilbert’s list, that of reconstructing a Fuchsian equation (1) from a collection of monodromy matrices associated with a set of points $a_{1},\dots,a_{N}\in\mathbb{C}$ [4, 2].
@@ 2.  Schlesinger system and superelliptic curves
Let $B^{(1)},B^{(2)},\ldots,B^{(N)}$ be $p\times p$ matrices depending on some complex parameters $a_{1},\ldots,a_{N}$ defined on a ball $D\in\mathbb{C}^{N}\setminus\{(a_{1},\ldots,a_{N})\,\,|\,\,a_{i}=a_{j}\,\text{for some }\,i\neq j\}$. The Schlesinger system for matrices $B^{(j)}$ has the form
@@
$\pi:X_{a}\,\backslash\,\pi^{-1}(\{\infty\})\rightarrow\hat{\Gamma}_{a}\,\backslash\,\{\infty\}$
@@
$\xi_{Q}(P)=\zeta(P)-\zeta(Q),$ $\mbox{if}\quad P\sim Q\quad\mbox{and $Q$ is a regular point.}$.
@@
[7] Let $X_{a}$ with $a=(a_{1},\dots,a_{N})\in D$ be a family of compact Riemann surfaces defined above corresponding to the family of algebraic curves (4). Let $m$ and $N$ be natural integers from (4) and let $n\in\mathbb{Z}$ be coprime to $m$. Let $\Omega_{i}^{j}(a)$ be the meromorphic differential defined by (6) on $X_{a}$. Then a set of upper triangular $p\times p$ matrices $B^{(i)}$, $i=1,...,N$ defined by the following three conditions provides a solution to the Schlesinger system (3):
@@
  • The eigenvalues of each matrix $B^{(i)}$, $i=1,...,N$, satisfy
@@
      $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s}\},\mathbb{C})$ if $m$, $N$ are not coprime, $s={\rm gcd}(m,N)$, and $n>0$.
@@
      $H_{1}(X_{a}\setminus\{P_{a_{1}},\ldots,P_{a_{N}}\},{\mathbb{C}})$ if $n<0\,.$
@@ 3.  Solutions of upper triangular Fuchsian systems
Let $p,\;n,\;m,\;N$ and $X_{a},a\in D$ be as in Theorem 1. Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ and let $\gamma_{j}$ be a closed contour on $X_{a}$ from Theorem 1 not passing through any point $P\in X_{a}$ for which $\zeta(P)=z$. More precisely, let $\gamma_{j}$ from Theorem 1 be seen as an element of
@@
    $H_{1}(X_{a}\setminus\{P\in X_{a}\,|\,\zeta(P)=z\},\mathbb{C})$ if $m$, $N$ are coprime and $n>0$,
  2. (a)
    $H_{1}(X_{a}\setminus(\{\infty_{1},\ldots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),\mathbb{C})$ if $n>0$ and $s={\rm gcd}(m,N)>1$,
@@
    $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),{\mathbb{C}})$ if $n<0$,
independent of small variations of $a\in D$. For such a choice of the contour $\gamma_{j}$, let $\kappa_{j}$ be a function of $z$ given by Definition 2. Let $B^{(i)},\;i=1,\dots,N,$ be upper triangular $p\times p$ matrices given by Theorem 1 with eigenvalues
@@
where $\theta_{i}\in\mathbb{R}$. Then, a fundamental upper triangular matrix solution $\Phi(z)$ to the Fuchsian system (1) with coefficients being the matrices $B^{(i)}$, $i=1,\ldots,N$, is given by
@@
(14) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
where $q\vdash(l-k)$ represents the partitions of the integer $l-k$, and $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
@@
The contour $\gamma_{j}$ for each $j=1,\dots,p-1$ in Theorem 4 is the same as the respective contour $\gamma_{j}$ in Theorem 1 used to construct the matrices $B^{(i)}$ serving as coefficients in (1). However, each set of contours $\gamma_{1},\dots,\gamma_{p-1}$ corresponding to a fixed solution $B^{(1)},\dots,B^{(N)}$ (7), (8) of the Schlesinger system and thus to a fixed Fuchsian system (1), gives rise to several different solutions of the Fuchsian system (corresponding to different initial conditions for (1)). This happens because, as stated in Theorem 4, the contours $\gamma_{j}$ are elements of the homology of the surface punctured at the points $P\in X_{a}$ for which $\zeta(P)=z.$ In other words, a given contour $\gamma_{j}$ in $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$ will give different values of the integral $\kappa_{j}$ (9) depending on the positions of the points of the surface satisfying $\zeta(P)=z_{0}$ relative to the contour $\gamma_{j}.$ Here we assume that initial value for $\Phi$ is given at $z=z_{0}$. Let us illustrate this with a simple example.
@@
    (19) $\kappa_{j}:=-\frac{m}{jn}\oint_{\gamma}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\underset{P_{z}^{(r)}}{\rm res}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}\Pi^{\frac{jn}{m}},$
    where $\epsilon_{r}$ are different $m$th roots of unity. Let us, for simplicity of notation, denote $\alpha_{j}:=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}$ so that $\kappa_{j}=\alpha_{j}\Pi^{\frac{jn}{m}}.$ Now, for the entries above the diagonal of the matrix $M$ given by (14), we have:
@@
    $\frac{\mathrm{d}}{\mathrm{d}z}D=D_{z}=\mathrm{diag}\left(\sum_{i=1}^{N}\frac{\beta_{1}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{1}^{(i)}},\ldots,\sum_{i=1}^{N}\frac{\beta_{p}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{p}^{(i)}}\right)$,
@@
    $(M_{z})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\sum_{r=1}^{l-k}\left(\frac{\kappa_{r}^{\sigma_{q}(r)-1}}{(\sigma_{q}(r)-1)!}\kappa_{r}^{\prime}\prod_{j=1,j\neq r}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\right)\quad&\text{when } $l>k$,\\ \quad 0\quad&\text{when } $l\leq k$,\end{dcases*}$
@@
    $D^{-1}=\mathrm{diag}\left(\prod_{j=1}^{N}(z-a_{j})^{-\beta_{1}^{(i)}},\ldots,\prod_{j=1}^{N}(z-a_{j})^{-\beta_{p}^{(i)}}\right)$, and
@@
    (21) $(M^{-1})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{(-\kappa_{j})^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$.\end{dcases*}$
@@
For $T=l-k$, this implies (22) with $l>k$ and thus proves that (21) defines the inverse of the matrix $M$ (14). $\Box$ We now have all the necessary tools to prove Theorem 4.
@@
The matrix $M_{z}M^{-1}$ is upper triangular as a product of upper triangular matrices. Its diagonal entries are zeros since this is the case for $M_{z}$, see Lemma 7. Let us consider the entries of $M_{z}M^{-1}$ lying above the main diagonal. Fix $k<l$ and set $j=l-k$. By Lemma 7, for $r\geqslant 1$, we have
@@
Let us now consider the contribution, denoted by $C(q,t),$ to (26) of a fixed partition $q\vdash j$ and of a fixed value of $t$. Introduce the vector $\sigma_{q}=(\sigma_{q}(1),\dots,\sigma(j))$, which defines the partition $q$ uniquely. We can rewrite the contribution of $q$ and $t$ in (26) as a sum over all decompositions of $\sigma_{q}$ into the sum of two vectors. Then, similarly to (23), this becomes the product of sums (we write $\sigma_{k}(i)$ for the $i$th component of the vector $\sigma_{k}$):
$C(q,t)=\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\end{subarray}}\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}=\sum_{\sigma_{1}+\sigma_{2}=\sigma_{q}}\sigma_{1}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{2}(i)}\binom{\sigma_{q}(i)}{\sigma_{1}(i)}$
@@
If $\sigma_{q}(i)\neq 0$, the sum over $x_{i}$ for $i\neq t$ equals $(1+x)^{\sigma_{q}(i)}$ evaluated at $x=-1$ and thus vanishes. If $\sigma_{q}(i)=0$, the sum over $x_{i}$ is equal to $1$. The sum over $x_{t}$, if $\sigma_{q}(t)>1$, can be rewritten as the derivative of $(1+x)^{\sigma_{q}(t)}$ evaluated at $-1$, therefore it is equal to $1$ if $\sigma_{q}(t)=1$ and 00 otherwise. Thus, $C(q,t)$ is non-zero if and only if $q=j$ and $t=j$, that is if $q$ is the partition of $j$ containing one part only, in which case $C(q,t)=1$. Using this result in (26), and noting that the only splitting of $j$ into the sum of two partitions that gives a non-zero contribution in (26) is $q_{1}=j,\;q_{2}=0$, we obtain
@@
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\!\sum_{h=k}^{l}M_{kh}(M^{-1})_{hl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{h=k}^{l}(h-k)M_{kh}(M^{-1})_{hl}$
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\delta_{kl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}.$
Comparing with our above derivation, we see that the second term can be rewritten using partitions similarly to (26), where $j=l-k$. The factor of $r$ corresponds to the sum of all the parts of the partition $q_{1}$. More precisely, using (14) for $M$ as well as formulas from (25) to (3), we obtain
@@
Given that $C(q,t)$ is non-zero only in the case of one term partition $q=j=l-k$ and $t=j,$ in which case $C(j,j)=1$, we obtain that quantity (28) is equal to $j\kappa_{j}$ and thus
$(MD_{z}D^{-1}M^{-1})_{kl}=\begin{cases}0,&l<k,\\ \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\ \displaystyle-\,\frac{(l-k)n}{m}\,\kappa_{\,l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
@@
$(\Phi_{z}\Phi^{-1})_{kl}=\begin{cases}0,&l<k,\\[3.0pt] \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\[4.0pt] \kappa_{l-k}^{\prime}-\displaystyle\frac{({l-k})n}{m}\,\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$

### v2

@@ 1. Introduction
Once a fundamental solution to (1) is found, one may obtain its monodromy matrices. The inverse problem, that of reconstructing a function $\Phi(z)$ from its monodromies, is referred to as a Riemann-Hilbert or inverse monodromy problem. Alternatively, it may be reformulated as a problem of finding two matrix functions, each of which is holomorphic either inside or outside of a closed contour passing through the points $a_{1},\dots,a_{N}$ in the complex plane, such that the two functions are related by the multiplication with a “jump” matrix on the contour. The monodromy matrices associated with the loops around $a_{i}$ may be combined to obtain this jump matrix [4, 17]. The term Riemann-Hilbert problem refers also to the 21st problem from Hilbert’s list, that of reconstructing a Fuchsian equation (1) from a collection of monodromy matrices associated with a set of points $a_{1},\dots,a_{N}\in\mathbb{C}$ [4, 2].
@@ 2. Schlesinger system and superelliptic curves
Let $B^{(1)},B^{(2)},\ldots,B^{(N)}$ be $p\times p$ matrices depending on some complex parameters $a_{1},\ldots,a_{N}$ defined on a ball $D\subset\mathbb{C}^{N}\setminus\{(a_{1},\ldots,a_{N})\,\,|\,\,a_{i}=a_{j}\,\text{for some }\,i\neq j\}$. The Schlesinger system for matrices $B^{(j)}$ has the form
@@
and we consider their family parametrized by $a=(a_{1},\ldots,a_{N})\in D$ with some fixed positive integers $m$ and $N$. The associated compact Riemann surfaces, denoted by $X_{a}$, are obtained by projectivizing and desingularizing the curves (4), see [18] and see also [8, 23] for application to superelliptic curves. Note that the algebraic curves (4) are nonsingular. The associated projective curves $\hat{\Gamma}_{a}\subset\mathbb{C}P^{2}$ may be singular at the points at infinity denoted collectively by $\{\infty\}$. By the theorem of resolution of singularities ([18], Chapter 7) there exists a holomorphic map $\pi:X_{a}\rightarrow\mathbb{C}P^{2}$ with $\pi(X_{a})=\hat{\Gamma}_{a}$ and such that the map
@@
$\xi_{Q}(P)=\zeta(P)-\zeta(Q),$ $\mbox{if}\quad P\sim Q\quad\mbox{and $Q$ is a regular point.}$
@@
[8] Let $X_{a}$ with $a=(a_{1},\dots,a_{N})\in D$ be a family of compact Riemann surfaces defined above corresponding to the family of algebraic curves (4). Let $m$ and $N$ be natural integers from (4) and let $n\in\mathbb{Z}$ be coprime to $m$. Let $\Omega_{i}^{j}(a)$ be the meromorphic differential defined by (6) on $X_{a}$. Then a set of upper triangular $p\times p$ matrices $B^{(i)}$, i=1,,Ni=1,...,N defined by the following three conditions provides a solution to the Schlesinger system (3):
@@
  • The eigenvalues of each matrix $B^{(i)}$, i=1,,Ni=1,...,N, satisfy
@@
      $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s}\},\mathbb{C})$ if $m$, NN are not coprime, $s={\rm gcd}(m,N)$, and $n>0$.
@@
      $H_{1}(X_{a}\setminus\{P_{a_{1}},\ldots,P_{a_{N}}\},{\mathbb{C}})$ if n<0.n<0\,.
@@ 3. Solutions of upper triangular Fuchsian systems
Let p,n,m,Np,\;n,\;m,\;N and $X_{a},a\in D$ be as in Theorem 1. Let $z_{0}\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ and let $\gamma_{1},\dots,\gamma_{p-1}$ be closed contours on $X_{a}$ from Theorem 1 not passing through any point $P\in X_{a}$ for which $\zeta(P)=z_{0}$. More precisely, let $\gamma_{j}$ from Theorem 1 be seen as an element of
@@
    $H_{1}(X_{a}\setminus\{P\in X_{a}\,|\,\zeta(P)=z_{0}\},\mathbb{C})$ if $m$, NN are coprime and $n>0$,
@@
    $H_{1}(X_{a}\setminus(\{\infty_{1},\ldots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z_{0}\}),\mathbb{C})$ if $n>0$ and $s={\rm gcd}(m,N)>1$,
  3. (c)
    $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z_{0}\}),{\mathbb{C}})$ if $n<0$,
independent of small variations of $a\in D$. Let $\mathcal{Z}\subset\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ be a small neighbourhood of $z_{0}$ not intersecting projections on the $\zeta$-sphere of the contours $\gamma_{1},\dots,\gamma_{p-1}$. For such a choice of the contour $\gamma_{j}$, let $\kappa_{j}$ be a function of $z\in\mathcal{Z}$ given by Definition 2. Let $B^{(i)},\;i=1,\dots,N,$ be upper triangular $p\times p$ matrices given by Theorem 1 with eigenvalues
@@
where $\theta_{i}\in\mathbb{R}$. Then, for $z\in\mathcal{Z}$ a fundamental upper triangular matrix solution $\Phi(z)$ to the Fuchsian system (1) with coefficients being the matrices $B^{(i)}$, $i=1,\ldots,N$, is given by
@@
[…]
where $q\vdash(l-k)$ represents the partitions of the integer lkl-k, and $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
@@
The contour $\gamma_{j}$ for each $j=1,\dots,p-1$ in Theorem 4 is the same as the respective contour $\gamma_{j}$ in Theorem 1 used to construct the matrices $B^{(i)}$ serving as coefficients in (1). However, each set of contours $\gamma_{1},\dots,\gamma_{p-1}$ corresponding to a fixed solution $B^{(1)},\dots,B^{(N)}$ (7), (8) of the Schlesinger system and thus to a fixed Fuchsian system (1), gives rise to several different solutions of the Fuchsian system (corresponding to different initial conditions for (1)). This happens because, as stated in Theorem 4, the contours $\gamma_{j}$ are elements of the homology of the surface punctured at the points $P\in X_{a}$ for which $\zeta(P)=z_{0}.$ In other words, a given contour $\gamma_{j}$ in $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$ will give different values of the integral $\kappa_{j}$ (9) depending on the positions of the points of the surface satisfying $\zeta(P)=z_{0}$ relative to the contour $\gamma_{j}.$ Here we assume that initial value for $\Phi$ is given at $z=z_{0}$. Let us illustrate this with a simple example.
@@
    (19) $\kappa_{j}:=-\frac{m}{jn}\oint_{\gamma}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\underset{P_{z}^{(r)}}{\rm res}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\epsilon_{r}\Pi^{\frac{jn}{m}},$
    where $\epsilon_{r}$ are different $m$th roots of unity. Let us, for simplicity of notation, denote $\alpha_{j}:=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\epsilon_{r}$ so that $\kappa_{j}=\alpha_{j}\Pi^{\frac{jn}{m}}.$ Now, for the entries above the diagonal of the matrix $M$ given by (14), we have:
@@
    $\frac{\mathrm{d}}{\mathrm{d}z}D=D_{z}=\mathrm{diag}\left(\sum_{i=1}^{N}\frac{\beta_{1}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{1}^{(j)}},\ldots,\sum_{i=1}^{N}\frac{\beta_{p}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{p}^{(j)}}\right)$,
@@
[…]
@@
    $D^{-1}=\mathrm{diag}\left(\prod_{j=1}^{N}(z-a_{j})^{-\beta_{1}^{(j)}},\ldots,\prod_{j=1}^{N}(z-a_{j})^{-\beta_{p}^{(j)}}\right)$, and
@@
[…]
@@
For T=lkT=l-k, this implies (22) with $l>k$ and thus proves that (21) defines the inverse of the matrix $M$ (14). $\Box$ We now have all the necessary tools to prove Theorem 4.
@@
The matrix $M_{z}M^{-1}$ is upper triangular as a product of upper triangular matrices. Its diagonal entries are zeros since this is the case for $M_{z}$, see Lemma 7. Let us consider the entries of $M_{z}M^{-1}$ lying above the main diagonal. Fix $k<l$ and set j=lkj=l-k. By Lemma 7, for $r\geqslant 1$, we have
@@
Let us now consider the contribution, denoted by $C(q,t),$ to (26) of a fixed partition $q\vdash j$ and of a fixed value of $t$. Introduce the vector $\sigma_{q}=(\sigma_{q}(1),\dots,\sigma_{q}(j))$, which defines the partition $q$ uniquely. We can rewrite the contribution of $q$ and $t$ in (26) as a sum over all decompositions of $\sigma_{q}$ into the sum of two vectors. Then, similarly to (23), this becomes the product of sums (we write $v_{k}(i)$ for the $i$th component of the vector $v_{k}$):
$C(q,t)=\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\end{subarray}}\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}=\sum_{v_{1}+v_{2}=\sigma_{q}}v_{1}(t)\;\prod_{i=1}^{j}(-1)^{v_{2}(i)}\binom{\sigma_{q}(i)}{v_{1}(i)}$
@@
If $\sigma_{q}(i)\neq 0$, the sum over $x_{i}$ for $i\neq t$ equals $(1+x)^{\sigma_{q}(i)}$ evaluated at x=1x=-1 and thus vanishes. If $\sigma_{q}(i)=0$, the sum over $x_{i}$ is equal to $1$. The sum over $x_{t}$, if $\sigma_{q}(t)>0$, can be rewritten as the derivative of $(1+x)^{\sigma_{q}(t)}$ evaluated at 1-1, therefore it is equal to $1$ if $\sigma_{q}(t)=1$ and 0 otherwise. Thus, $C(q,t)$ is non-zero if and only if $q=j$ and $t=j$, that is if $q$ is the partition of $j$ containing one part only, in which case $C(q,t)=1$. Using this result in (26), and noting that the only splitting of $j$ into the sum of two partitions that gives a non-zero contribution in (26) is $q_{1}=j,\;q_{2}=0$, we obtain
@@
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}}\!\sum_{h=k}^{l}M_{kh}(M^{-1})_{hl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{h=k}^{l}(h-k)M_{kh}(M^{-1})_{hl}$
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}}\delta_{kl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}.$
Comparing with our above derivation, we see that the second term can be rewritten using partitions similarly to (26), where j=lkj=l-k. The factor of $r$ corresponds to the sum of all the parts of the partition $q_{1}$, that is $r=\sum_{t=1}^{j}t\,\sigma_{q_{1}}(t)$. More precisely, using (14) for $M$ as well as formulas from (25) to (3), we obtain
@@
Given that $C(q,t)$ is non-zero only in the case of one term partition q=j=lkq=j=l-k and $t=j,$ in which case $C(j,j)=1$, we obtain that quantity (28) is equal to $j\kappa_{j}$ and thus
$(MD_{z}D^{-1}M^{-1})_{kl}=\begin{cases}0,&l<k,\\ \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}},&l=k,\\ \displaystyle-\,\frac{(l-k)n}{m}\,\kappa_{\,l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
@@
$(\Phi_{z}\Phi^{-1})_{kl}=\begin{cases}0,&l<k,\\[3.0pt] \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}},&l=k,\\[4.0pt] \kappa_{l-k}^{\prime}-\displaystyle\frac{({l-k})n}{m}\,\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$